OpenStudy (jhannybean):

Write the equation of a plane : \[\vec r(u,v) =2\sin(u)\hat i + 3\cos(u)\hat j +v\hat k\] \(0\le v\le 2\)

OpenStudy (anonymous):

some vector stoof

OpenStudy (dan815):

r_u X r_v

OpenStudy (jhannybean):

\[P_0 = (2,3,0)\]

OpenStudy (dan815):

will give yu normal line <a,b,c> d=ax+by+cz

OpenStudy (dan815):

plug po in and see what d has to be

OpenStudy (anonymous):

I can do this but I can't really visualize this haha.

OpenStudy (dan815):

basicalllyy

OpenStudy (jhannybean):

Uhh... vector a.... or \(\vec r_u = <\cos(u) ~ , ~ -\sin(u)~,~0>\)

OpenStudy (dan815):

hm

OpenStudy (jhannybean):

\[\vec b = <0,0,1>\]

ganeshie8 (ganeshie8):

you're trying to find the tangent plane of given surface at given point ?

OpenStudy (jhannybean):

Im so lazy rn.

OpenStudy (jhannybean):

Yeah.

OpenStudy (dan815):

i know what we find can be the tangent pane, but can u say this whole thing iss a plane

ganeshie8 (ganeshie8):

that whole thing looks like a cylinder to me

OpenStudy (anonymous):

Dan this is the first and last time you'll hear this, can you draw it out please?

ganeshie8 (ganeshie8):

it is aligning in z axis

OpenStudy (dan815):

its an elliptic cylinder in u,v

OpenStudy (jhannybean):

\[a \times b =\]